A grocer mixes 5 lb of egg noodles costing .80/lb with 2 lb of spinach noodles costing $1.50/lb. What will the cost per pound of the mixture be?
Answer by checkley79(3341)

5*.80+2*1.50=7x
4.00+3.00=7x
7x=7.00
x=7.00/7
x=$1.00 cost for the mix.